OpenStudy (anonymous):

If g(x) = 3x + 5, what is g-1(x)?

hero (hero):

1. Let y = g(x). After doing so, the equation becomes y = 3x + 5 2. Swap x and y x = 3y + 5 3. Isolate y: \[\frac{x - 5}{3} = y\] 4. Rewrite as \[y = \frac{x - 5}{3}\] 5. Replace y with \(g^{-1}(x)\): \[g^{-1}(x) = \frac{x - 5}{3}\]

OpenStudy (anonymous):

Oh that's what the -1 means, that's right. Thanks, Hero.

hero (hero):

-1 in the exponent always means "inverse"

hero (hero):

You pronounce \(g^{-1}(x)\) as "g inverse of x"...as I'm sure you already know.

OpenStudy (anonymous):

Can you show me the steps between step 2 and step 3?

OpenStudy (anonymous):

Nevermind, I get it. XD

hero (hero):

Okay. I try to make my steps clear.

hero (hero):

I know some steps may be unclear. It is good that you try to understand then ask questions if you don't. Most students would just stare at it without speaking. Then leave assuming that I will only confuse them more.

hero (hero):

I assure you that I have the ability to make things that are confusing clearer and not more confusing.

OpenStudy (anonymous):

Oh no doubt. It just took me a moment to piece together how to solve for y. That's all. :)

OpenStudy (zzr0ck3r):

listen to him^^^:)

hero (hero):

Thanks for the support @zzr0ck3r. You got back pretty quick bro.

OpenStudy (anonymous):

One more question, What is the description of f(x) = -3 - x^6? The exponent is 6 so it's an even graph, and the 3 is a negative so it curves to the left?

OpenStudy (anonymous):

Err *decreases to the left

hero (hero):

When figuring this out, you want to begin with the lead term, which is x^6. So what you do is say that f(x) is the graph of x^6 flipped over the horizontal axis and decreased by 3.

hero (hero):

If you graph f(x), you will see that this is true.

OpenStudy (anonymous):

But it is an even function?

hero (hero):

You can also mention that it is an even function. You figure that out by replacing x with -x and then simplifying: f(-x) = -((-x)^6 - 3 f(-x) = - x^6 -3 You know it's an even function when f(-x) = f(x)

OpenStudy (anonymous):

So you don't focus on the exponent when determining this?

hero (hero):

When determining it, you focus on the definition of an even function. That is f(x) = f(-x). That's how you determine it. Focusing on the exponent is like making assumptions. Many people make assumptions with mathematics and that's what leads to mistakes. What if you had: f(x) = x^2 - x Now let's replace that with f(-x) and see what we get: f(-x) = (-x)^2 -(-x) f(-x) = x^2 + x We did not arrive back at the original expression even though we had an even exponent.

OpenStudy (anonymous):

And by not arriving back at the original equation, it's an odd function?

hero (hero):

It's only an odd function if f(-x) = -f(x) -f(x) = -(x^2 - x) = -x^2 + x So it's not an odd function either.

hero (hero):

It is neither even nor odd. So as you can see, we have to use the definition of even and odd function to determine what kind of function it is. Making assumptions will not work. Mathematics is solved by using definitions, properties, postulates, and theorems. You're allowed to use assumptions when proving theorems and such, but with computation, you have to be careful.

hero (hero):

When you make assumptions in mathematics, they have to be provable. Otherwise they are of no good use.

OpenStudy (anonymous):

Got it, I understand now. I don't even remember learning how to tell whether it was even or odd. But I'm glad this got cleared up before my final. Thanks for the help!

hero (hero):

yw. @zzr0ck3r did you want to add anything?

hero (hero):

Okay, well, I'm out. ttyl

OpenStudy (anonymous):

Goodnight. Thank you again :)

hero (hero):

Bye

OpenStudy (zzr0ck3r):

@hero not at all, I was just watching and eating and agree with you:)

hero (hero):

I figured you must have been eating.